OpenStudy (anonymous):

What are the intercepts for the line 3x - 6y = 6?

OpenStudy (anonymous):

x-intercept (y=0) y-intercept (x=0) 3x - 6(0) = 6 3(0) - 6y = 6 3x = 6 -6y = 6 x = 6/3 y = 6/-6 x = 2 y = -1 (2,0) (0,-1)

OpenStudy (anonymous):

Thank you